Operations Report April 16th

Crew 245 Operations Report 16-04-2021

SOL: 6

Name of person filing report: Shravan Hariharan

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Generator: Working nominally

Hours run: 6.5

From what time last night: 10:30 pm, SOC 53%

To what time this morning: 5:00 am, SOC 100%

List any additional daytime hours when the generator was run: N/A

Solar— SOC 69% (Before generator is run at night),

Diesel Reading – 50%

Station Propane Reading – 73%

Water (loft tank): 25 gallons

Water Meter: 150396.0 units

Water (static tank): 300 gallons

Static to Loft Pump used – yes

Water in GreenHab: 0 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: No

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

Spirit rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: yes, in town

Opportunity rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: no, in town

Curiosity rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: yes, in town

Perseverance rover used: Yes

Hours: 202.1

Beginning charge: 100%

Ending charge: 99%

Currently charging: Yes

Notes on rovers: Nothing to report

ATV’s Used: 350.1

Reason for use: EVA

Oil Added? No

ATV Fuel Used: 0.5 gallon (estimated)

# Hours the ATVs were used today: 0.75 (but 0.5 hours of active transit)

Notes on ATVs: Mission Director will monitor and refuel ATVs for duration of mission

HabCar used and why, where?: No

CrewCar used and why, where?: No

General notes and comments: Nothing to report

Summary of internet: Internet is working nominally. Crew did not consume allotted bandwidth last night so there was an additional bandwidth bonus available today.

Summary of suits and radios: All radios nominal, all suits currently charging. Suits 7, 9, and 11 used on EVA today and functioned nominally.

Summary of Hab operations: Nothing to report besides nominal crew activities.

Summary of GreenHab operations: Being used to host Crew Botanist botany experiment. Crew Scientist moved hydroponics experiment from GreenHab to lower deck of Hab, as intense heat and sunlight in GreenHab was causing cuttings to dry up.

Summary of ScienceDome operations: Used by HSO for mini-botany experiment (Space Farmer Dream Kit), with assistance from the Crew Botanist.

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report.


Shravan Hariharan, Crew Engineer